Barnett Products manufactures three types of remote-control devices: Economy, Standard, and Deluxe. The company, which uses activity-based costing, has identified five activities (and related cost drivers). Each activity, its budgeted cost, and related cost driver is identified below.

Activity Cost Cost Driver
Material handling $ 225,000 Number of parts
Material insertion 2,475,000 Number of parts
Automated machinery 840,000 Machine hours
Finishing 170,000 Direct labor hours
Packaging 170,000 Orders shipped
Total $ 3,880,000


The following information pertains to the three product lines for next year:

Economy Standard Deluxe
Units to be produced 10,000 5,000 2,000
Orders to be shipped 1,000 500 200
Number of parts per unit 10 15 25
Machine hours per unit 1 3 5
Labor hours per unit 2 2 2


What is Barnett’s pool rate for the finishing activity?

Multiple Choice

  • $5.00 per labor hour.

  • $5.00 per machine hour.

  • $5.00 per unit.

  • $7.50 per unit.

  • None of the answers is correct.
Answer #1

direct labor hours=(Units to be produced*Labor hours per unit)

direct labor hours for:

Economy=(10,000*2)=20,000

Standard=(5000*2)=10,000

Deluxe=(2000*2)=4000

Total direct labor hours=(20,000+10,000+4000)=34000

pool rate for the finishing activity=Respective cost/Respective activity

=170,000/34000

=$5 per labor hour